# SLIM: Sim-to-Real Legged Instructive Manipulation via Long-Horizon Visuomotor Learning

### v2

@@ III-C Teacher-Student Training
Given the challenges of an enormous exploration space implied by long-horizon tasks, we adopt a teacher-student learning framework  [11, 30, 60] for better high-level training efficiency. The overall structure of the framework is shown in the second and third blocks of the right side of Fig. 2. As shown, the teacher is trained purely with RL from privileged, structured, and low-dimensional environment data $(k,\mathbf{s_{\textrm{p}}},\mathbf{s_{\textrm{r}}})$ that can only be obtained from the simulator. After a high performing teacher policy is learned, it is frozen and used to guide the student’s policy learning via a distillation loss. As the student will eventually be deployed in the real world, its inputs are consistent with what our robot system has. Besides policy distillation, we also use an RL objective to allow the student’s behavior being shaped by the same set of task rewards used by the teacher. Later, we will show improvements of this distillation-guided RL setting, compared to pure policy distillation (Section VIII-B Distillation Only baseline.). In the next two sections, we will provide more technical details of the teacher and student.
@@
IV The Teacher: Progressive Policy Expansion for Long-Horizon Task Learning
IV-A Long-Horizon Task Learning Challenges
Since we are targeting at solving long-horizon tasks via RL, we need an effective approach to train a policy to solve long-horizon tasks successfully. There are are few key challenges that must be addressed for solving the long horizon tasks:
@@ IV-A Long-Horizon Task Learning Challenges
    Continual exploration: When learning a long-horizon task via RL, there may be a number of intermediate milestones that must be sequentially achieved. Therefore, the teacher must be able to keep exploring at its frontier despite solving an intermediate milestone stage, in pursuit of solving the full task. Without the ability of continual exploration, it is more likely to converge to a sub-optimal solution that solves the task only up to some intermediate milestones, therefore cannot solve the full task completely.
@@
    Loss of plasticity and catastrophic forgetting: As the teacher progresses into a task, it needs to properly cope with both the loss of capacity/plasticity [32, 33] of the network due to the evolving nature of the problem as well as the data imbalance during long-horizon RL task exploration and training, without impacting skills that have already been acquired (catastrophic forgetting).
IV-B Progressive Policy Expansion
Given the challenges listed above, we build upon the idea of Policy Expansion (PEX) [58] originally used in the context of offline-to-online RL. Intuitively, PEX divides the full learning process into stages (e.g. offline and online stages in [58]) and uses dedicated networks for learning within each stage. This way, the trainer allocates new network capacity for continual exploration and learning of new skills without affecting any existing skill (catastrophic forgetting).
Here we generalize this idea for addressing the challenges in long-horizon task learning as mentioned in Section IV-A, by extending it to multiple stages. More concretely, we partition the full long horizon task into a sequence of stages (subtasks), and then perform a policy expansion operation for each new encountered subtask. By doing this, it can effectively combat the challenges in long-horizon task learning, keep exploring beyond the currently solved subtasks to solve the subsequent subtasks, circumvent the loss of plasticity by leveraging the newly allocated policy network  [^1] We use the term policy network to represent all the networks that are required for learning a policy. For example, in the context of Actor-Critic formulation, it encompasses both the actor and critic networks. for learning, and avoid catastrophic forgetting by retaining the policy networks for previous subtasks. Given these progressive multi-stage expansions, we dub this strategy Progressive Policy Expansion.
Formally, we represent the full long-horizon task $\mathcal{T}$ as a composition of a set of subtasks
$\mathcal{T}=\mathcal{C}(\{\tau^{k}\}_{k=1}^{K})$ (1)
where $\mathcal{C}$ is a composition operator, assembling the set of subtasks $\{\tau^{k}\}$ into the full task $\mathcal{T}$. $K$ denotes the maximum number of subtasks. We also define a task decomposional operator as follows, mapping an input state to the corresponding subtask index:
$k=\mathcal{D}(\mathbf{s})\quad k\in\{1,2,\cdots,K\},$ (2)
where $\mathbf{s}=[\mathbf{s}_{\textrm{r}},\mathbf{s}_{\textrm{p}}]$. We use $\pi_{\rm tea}^{k}(\mathbf{a}|\mathbf{s})$ to denote the teacher sub-network that is responsible for solving its corresponding subtask $\tau^{k}$. When there is no confusion, we omit the subscript of the teacher network $\pi_{\rm tea}^{k}$ as $\pi^{k}$. Figure 3 provides graphical illustrations of the teacher sub-network structures. Each sub-network (Figure 3 left) takes low-dimensional proprioceptive and privileged observation as input and is responsible for learning to solve the corresponding subtask.
@@ IV-B Progressive Policy Expansion
Figure 3: Teacher Policy Network Structure. On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP. In the case of actor-critic learning, both critic and actor networks have the same architecture, with the difference in that the actor network outputs an action distribution while the critic outputs a scalar value. On the right, the full teacher network $\Pi$ is composed of multiple structurally identical networks $\{\pi^{k}\}_{k=1}^{K}$. Given an input id $k$, the $k$-th component network is activated in the computation, i.e. $\Pi[k]\equiv\pi^{k}$.
We can then solve the long horizon task $\mathcal{T}$ via Progressive PEX as follows. We initiate the exploration and learning with a single policy network $\Pi=\{\pi^{1}\}$. This policy network is responsible for learning to solve the initial subtask. Whenever a new subtask is encountered, a policy expansion operation is conducted, adding a new policy into the policy set $\Pi=\{\pi^{1},\pi^{2}\}$. By doing this progressively for the full task, we get
@@
as illustrated in Figure 3 right. A forward pass through $\Pi$ is carried out via the proper sub-policy by indexing the policy set with the task decompositional operator $\mathcal{D}$ conditioned on the current state
$\Pi(\mathbf{a}_{\text{hi}}|\mathcal{D},\mathbf{s})$ $=\pi^{\mathcal{D}(\mathbf{s})}(\mathbf{a}_{\text{hi}}|\mathbf{s}).$ (4)
There are many possible ways to implement the task decompositional operator $\mathcal{D}$. In our case, since it is used in teacher policy learning, we can leverage the fact that privileged observation can be used by the teacher policy and implement it as part of the privileged observation provided by the simulator, e.g., providing the subtask index $k$ directly in the privileged observation. Based this implementation, the training of $\Pi$ is closely related to multi-task RL [56, 46], where we sample a batch of multi-task-transitions (transitions containing the task index $k$) from the replay buffer and use each transition in the batch for training the sub-network associated with its corresponding task index. In this work, we train the teacher policy using a multi-task variant of SAC algorithm [19, 56, 46],
@@ V The Student: Policy Distillation Guided RL
The student needs to perceive the surrounding environment through RGB images. There is no input regarding which subtask the student is currently solving, and it has to infer this information from the inputs. Because of this, the student learns only a single policy across the entire long-horizon task (no sub-task dedicated policy).
To tackle these challenges, we train the student by distilling the teacher’s multiple sub-task policies into a single task policy $\pi_{\rm stu}$. However, distillation will only give us a policy that works fine but not quite well. It could lead to the state drifting or out-of-distribution (OOD) issues when the policy is rolled out for a long horizon. There is also always an observation gap between the teacher and the student, which sometimes creates a difficulty in minimizing the distillation loss. Moreover, due to the differences in model architectures and inductive biases, the student might need a different way of mastering certain skills. Finally, distillation makes the student’s performance upper bounded by the teacher’s. Given these considerations, we choose to boost the student’s policy with RL under the same set of task rewards used by the teacher (Fig. 4).
We train the student by modifying SAC [19] to incorporate the distillation loss properly. First, we use a mixed rollout strategy to generate replay data. When rolling out a new episode, with a probability of $\beta$ we will sample from the student policy, or otherwise from the teacher policy. On one hand, we wish to use the teacher to efficiently generate high-performing transitions along the long horizon. On the other hand, we want to keep exploring with the student’s own policy. Second, following SACLite [55] we remove the entropy reward from policy evaluation. Then for policy improvement, we replace the entropy term with the distillation loss, and assign a fixed weight $\alpha$ to it:
$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{o},\mathbf{s}_{r},\mathbf{s}_{p},l,k)\sim\mathcal{D}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)}Q\big{(}(\mathbf{o},\mathbf{s}_{r},l),\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha KL[\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{r},\mathbf{s}_{p})||\pi_{\rm stu}(\cdot|\mathbf{o},\mathbf{s}_{r},l)]\Big{]}.\\ \end{array}$ (5)
To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform a policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion $\sigma$ (e.g., std. for Gaussian). This makes our distillation loss the best of two worlds: imitation learning and entropy regularization.
@@
Figure 4: A close look at the student’s architecture. Red dashed lines represent data flows that only exist for training in simulation. Solid lines denote the inference process of the student policy. For deployment, the three inputs are replaced with real-world observations.
When designing the student’s representation model, we specifically take vision sim2real gap reduction into consideration, because once after the student is trained, it will be directly deployed in real without any finetuning. While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck to reduce the vision sim2real gap (Fig. 4). The idea is to let the policy depend on as minimal visual information as possible, namely, to limit the complexity of its input vision space. This potentially leads to a better generalization ability. While there might be many choices for the vision bottleneck, from discrete latent codes [20] to canonical images [25], we choose to use a pair of segmentation and depth maps. There are mainly two reasons for this choice. First, we believe that they contain the minimal visual information required to achieve our task, while without incurring much task information loss. Second, our simulator already provides groundtruth segmentation and depth maps, as high-quality supervision for learning such a vision bottleneck. Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs. Note that only the robot and task-related objects appearing in the instruction will be predicted in the segmentation mask. All other irrelevant object pixels will be classified as “background”. After this, the segmentation and depth maps will be fed to the downstream policy for generating actions. Using a segmentation mask for selecting interesting objects is similarly adopted by prior works [30, 53].
@@ VI Task Implementation In Simulation
Figure 5: Task Decomposition as Privileged Information. Given a task, we decompose it into a sequential series of primitive subtasks. As shown above, the task of finding an object A, grasping it, and then dropping into an object B can be decomposed into a series of subtasks. In simulation, the subtask boundaries and transitions are well-defined, serving as a form of privileged information to the teacher. The teacher can leverage this information, and use a separate policy to learn each subtask, thereby conquering the long horizon task. This expertise is then distilled into the student, which is conditioned on both visual camera feed and language commands.
We use a long-horizon pick-and-place task as a concrete example in this work (c.f. Figure 2 left for an illustration of the task), and we will use this task as a concrete example for illustrating the task decomposition as well. Given a language instruction (e.g., “Drop the blue cube into the green basket.”), the robot needs to locate the target object, picking it up and then drop it into the target basket.
As mentioned previously in Section IV, SLIM employs a task decompositional approach paired with progressive PEX for training a teacher policy to solve the long horizon task. More concretely, for the long-horizon task under consideration, it can be naturally broken down into the following sequential steps (c.f. Figure 5):
@@
    Search: the robot should first search for the target object (the green cube) for picking up (a subtask), based on the RGB images from the gripper mounted ego-centric camera;
@@
    MoveTo: after finding the target object, the robot moves to the object until it is within the reach of the arm;
@@
    Grasp: then the robot moves the gripper towards the target object for picking it up. During the grasping process, the robot base should ideally move less compared to the MoveTo subtask to facilitate accurate grasping;
@@
    SearchWObj: after picking up the target object, the robot should keep it in the gripper and then search for the target basket; within this stage, there should be some coordination between arm and locomotion movement to facilitate search while avoid the cube to be dropped.
@@
    MoveToWObj: after finding the target basket, the robot should move towards it while keeping the object in gripper, until the target basket is within a range that is appropriate for the robot to reach for dropping the object in its gripper;
@@
    MoveGripperToWObj: the robot can now move the gripper towards the target basket until reaching an appropriate position on top of the basket for dropping;
@@
    DropInto: the robot releases its gripper and drops the cube into the basket.
In addition, we can further append an auxiliary Idle subtask, which encourage the robot to adjust its arm back to neutral position after dropping the object into the basket, facilitating subsequent deployment. This is an optional subtask that is not required for task completion. We include it in our task implementation to make the final policy behavior safe and more predictable after completing the full task. When evaluating the real world performance (e.g. Section VIII-D), we do not include this auxiliary subtask into the full task metric. We also exclude it whenever we discuss task-relevant decompositions since it is not a task related subtask.
Formally, we instantiate the task compositional operator as
$\mathcal{T}=\mathcal{C}(\{\tau^{k}\}_{k=1}^{K})\triangleq\tau^{1}\rightarrow\tau^{2}\rightarrow\cdots\rightarrow\tau^{K}$ (6)
where $\tau^{i}\!\rightarrow\!\tau^{j}$ denotes that the two subtasks $\tau^{j}$ and $\tau^{j}$ are two adjacent subtasks and $\tau^{j}$ follows subtask $\tau^{i}$ in a sequential manner. In our case, $K\!=\!7$ and $\{\tau^{1},\tau^{2},\cdots,\tau^{7}\}$ corresponds to {Search, MoveTo, Grasp, SearchWObj, MoveToWObj, MoveGripperToWObj, DropInto}.
Because of the decomposition, we can implement reward functions for each subtask, making it modular and easier for implementation. From some tasks, we simply use the sparse subtask success reward (e.g. Search). We incorporated delta-distance based shaping reward for tasks that requires to move the robot (or gripper) close to a target region (e.g. MoveTo, MoveToWObj, MoveGripperToWObj, Grasp, DropInto), in addition to the sparse subtask reward. For SearchWObj and MoveToWObj, we also incorporated an arm retract reward, which is another instance of delta-distance based shaping reward, where the distance is measured between the current arm-joint pose (excluding gripper) to a target neutral arm pose (c.f. Figure A.13). Intuitively, the arm retract reward will encourage the robot’s arm to raise up after grasping the object and move close to its neutral position. Its impact will be investigated in Section VIII-B (No Arm Retract baseline).
An additional benefit of task decomposition is that we have the option to incorporate behavior priors in a flexible way into the policy. More details on the subtasks, rewards and behavior priors are provided in Appendix A-A.

### v3

@@ III-C Teacher-Student For High-Level Training
With the low level policy taking over the responsibilities of legged locomotion control, high level can focus on task dependent decision making. However, the high level policy needs to process high dimensional visual and language input, and complete the long-horizon task. This presents an enormous space for the policy to explore, and a challenge for efficient RL training. We adopt a teacher-student learning framework  [12, 33, 66] for efficient high-level training. The overall structure of the framework is shown in the middle and bottom right sections of Fig. 2. As shown, the teacher is trained purely with RL from privileged, structured, and low-dimensional inputs $(k,\mathbf{s_{\textrm{p}}},\mathbf{s_{\textrm{r}}})$ some of which can be obtained from the simulator but not easily in the real world. When a teacher policy is successfully learned, it is frozen and used to guide the student via behavior distillation. As the student will eventually be deployed in the real world, its inputs no longer contain privileged information. Besides policy distillation, we also use an RL objective to allow the student’s behavior to be shaped by the same set of task rewards used by the teacher, for a potentially more effective policy than using distillation alone (details in Section VIII-B, Distillation Only baseline). In the two sections below, we explain the teacher and the student policies in more detail.
@@
IV The Teacher: Long-Horizon RL with Task Decomposition and Policy Expansion
The Teacher works with privileged object state information, thus, offloading the complexities of visual representation learning to the student, and only needs to focus on the following challenges of long-horizon task learning,
@@ IV The Teacher: Long-Horizon RL with Task Decomposition and Policy Expansion
    Continual exploration: For long-horizon tasks, there may be a number of intermediate milestones that must be sequentially achieved. Therefore, even after reaching an intermediate milestone, the teacher must continue exploring new frontiers in order to ultimately solve the entire task. Without carefully encouraging continual exploration, the agent can stop exploration early and settle on a suboptimal solution.
@@
    Loss of capacity and catastrophic forgetting: As the teacher progresses in training, it needs to cope with both the loss of capacity issue [35, 36], which hinders the network from continual learning, and the catastrophic forgetting issue, which could destroy skills that have already been acquired during the learning of new tasks.
Addressing these challenges are key for successful teacher policy learning. We propose to use task decomposition and further integrate with policy expansion (PEX) [64] to address these two challenges.
We first decompose the long horizon task $\mathcal{T}$ into $K$ subtasks $\{\tau^{k}\}_{k=1}^{K}$ with shorter-horizons. To incorporate the task decompositional structure in learning, we leverage the fact that privileged information is accessible to the Teacher and thus include the subtask index $k$ as part of the privileged observation. Subtask successes keep pushing the agent to learn new subtasks one after another, addressing the exploration issue.
Secondly, we solve the loss of plasticity and forgetting in long horizon task learning via a progressive application of the PEX scheme [64]. Originally designed for two-stage offline-to-online training, PEX uses dedicated networks for learning within each stage. This way, the trainer creates new network instances for continual exploration and learning of new skills without affecting any skills acquired in the previous stage. Here we apply this idea progressively to the decomposed long-horizon task with multiple subtasks, where the stages corresponds to the subtasks. We dub this strategy as Progressive PEX.
Because of policy expansion at each stage transition in Progressive PEX, it addresses the two challenges brought by the long-horizon task learning naturally: 1) it can achieve multi-stage continual exploration and the full ability of exploration is always ensured when entering a new stage; 2) since the policy for solving one subtask is encapsuled in a dedicated network, it mitigates the catastrophic forgetting issue for the already learned subtask policies. For a newly encountered subtask, the dedicated network that will be newly allocated addresses the issue of lost plasticity.
@@
Figure 3: Teacher Policy Network Structure. The teacher network is a set of identical submodules $\{\pi^{k}\}_{k=1}^{K}$ gated by the subtask id $k$ (right side of figure). Each submodule $\pi^{k}$ is only trained with the experience when the agent is working on subtask $k$, taking privileged and proprioceptive input, flattening and concatenating them before passing it through an MLP (left side of figure). On the left, each individual teacher policy $\pi^{k}$ takes a set of privileged and proprioceptive information as input, flattening and concatenating them before passing it through an MLP. On the right, the full teacher network $\Pi$ is composed of multiple structurally identical networks $\{\pi^{k}\}_{k=1}^{K}$. Given an input id $k$, the $k$-th component network is activated in the computation, i.e., $\Pi[k]\equiv\pi^{k}$.
We use $\pi_{\rm tea}^{k}(\mathbf{a}_{\text{hi}}|\mathbf{s})$ to denote the teacher sub-network that is responsible for solving its corresponding subtask $\tau^{k}$. When there is no confusion, we omit the subscript of the teacher network $\pi_{\rm tea}^{k}$ as $\pi^{k}$. The left side of Fig. 3 provides a graphical illustration of the teacher sub-network structures. Each sub-network takes low-dimensional proprioceptive and privileged observation as input and is responsible for learning to solve the corresponding subtask.
Progressive PEX works as follows. We initiate the exploration and learning with a single policy network [^1] We use the term policy network to represent all the networks that are required for learning a policy. For example, in the context of Actor-Critic formulation, it encompasses both the actor and critic networks. $\Pi=\{\pi^{1}\}$ that is responsible for learning to solve the initial subtask. Whenever a new subtask is encountered, a policy expansion operation is conducted, adding a new policy into the policy set, i.e., $\Pi=\{\pi^{1},\pi^{2}\}$. By doing this progressively for the all the $K$ subtasks, we get
@@
as illustrated in the right side of Fig. 3.
Based on this implementation, the training of $\Pi$ is closely related to multi-task RL [62, 49]. For each training iteration, we get a batch of samples containing the subtask indices from the replay buffer and use each sample in the batch for training the sub-network associated with the subtask index of that sample. In this work, we train the teacher policy using a multi-task variant of SAC algorithm [21, 62, 49]. The performance comparison between the method using a standard network structure and the Progressive PEX approach on solving long-horizon tasks is provided in Appendix J-B, which clearly shows Progressive PEX has a much stronger ability in long-horizon task learning.
@@ V The Student: Policy Distillation Guided RL
The student needs to perceive the surrounding environment using RGB, motor sensors and language instruction. Because the student needs to be deployed in the real world, it cannot use any privileged information e.g. which subtask it needs to solve. Therefore, the student has to learn a single policy over the entire long-horizon task.
To tackle these challenges, we train the student by distilling the teacher’s behavior into a single task policy $\pi_{\rm stu}$. However, distillation on its own is likely to produce a policy that works reasonably well but it could lead to state drifting / out-of-distribution (OOD) issues when the policy is rolled out over a long horizon. There is also always an observation gap between the teacher and student, which may create difficulty in minimizing the distillation loss. Moreover, due to differences in model architectures and inductive biases, the student may need some adaptation of the teacher’s skills. Given these considerations, we choose to boost the student’s policy with RL under the same set of task rewards used to train the teacher (Fig. 4).
We train the student by modifying SAC [21] to incorporate the distillation loss properly. First, we use a mixed rollout strategy to generate replay data. When rolling out a new episode, with a probability of $\beta$, we will sample actions from the student policy, otherwise from the teacher policy. On the one hand, we wish to use the teacher to efficiently generate high-performing transitions along the long horizon. On the other hand, we want to keep exploring with the student’s own policy. Second, following SACLite [61], we remove the entropy reward from policy evaluation. Then for policy improvement, we replace the entropy term with the distillation loss and assign a fixed weight $\alpha$ to it:
$\begin{array}[]{l}\displaystyle\max_{\pi_{\rm stu}}\displaystyle\mathbb{E}_{(\mathbf{s}_{\textrm{stu}},\mathbf{s}_{\textrm{p}},k)\sim\mathcal{D}_{\text{replay}}}\Big{[}\mathbb{E}_{\mathbf{a}_{\text{hi}}\sim\pi_{\rm stu}(\cdot|\mathbf{s}_{\textrm{stu}})}Q\big{(}\mathbf{s}_{\textrm{stu}},\mathbf{a}_{\text{hi}}\big{)}\\ \ \ \ \ \ \ \ -\alpha{\rm KL}\!\left(\pi_{\rm tea}^{k}(\cdot|\mathbf{s}_{\textrm{r}},\mathbf{s}_{\textrm{p}})\,||\,\pi_{\rm stu}(\cdot|\mathbf{s}_{\textrm{stu}})\right)\Big{]},\\ \end{array}$ (2)
where $\mathbf{s}_{\textrm{stu}}=[\mathbf{o},\mathbf{s}_{\textrm{r}},\mathbf{l}]$. To ensure that the KL term encourages enough randomness and exploration for $\pi_{\rm stu}$, we perform policy surgery on $\pi_{\rm tea}$ where we keep the action distribution mode unchanged but assign a fixed modal dispersion $\sigma$ (e.g., std. for Gaussian). This results in a distillation loss that elegantly achieves two goals: imitation learning and entropy regularization.
@@
Figure 4: Student Architecture Overview. Red dashed lines represent data flows that only exist in simulation training. Solid lines denote the inference process of the student policy.
When designing the student’s representation model, we specifically take visual sim-to-real gap reduction into consideration, because once the student is trained, it will be directly deployed in real without any finetuning. While the student could have a naive representation model that simply fuses multimodal inputs and generates a latent encoding for its policy to use, we choose to add an intermediate visual information bottleneck (Fig. 4) to reduce the visual sim-to-real gap while achieving better interpretability.
The idea is to let the policy depend on as minimal visual information as possible, namely, to limit the complexity of its input vision space. This can leads to potentially better generalization. While there might be many choices for the vision bottleneck, from discrete latent codes [22] to canonical images [28], we choose to use a pair of segmentation and depth maps. There are mainly three reasons for this choice. First, we believe that they contain the minimal visual information required to achieve our task without incurring much task-relevant information loss. Second, ground-truth segmentation and depth maps can be efficiently obtained from the simulator for high-quality supervised learning of such a vision bottleneck. Thus when training the student, we add a supervision loss for predicting the segmentation mask and depth map from its inputs. Note that only the robot and task-related objects appearing in the instruction will be predicted in the segmentation mask. All other irrelevant object pixels will be classified as “background”. After this, the segmentation and depth maps will be fed to the downstream policy for generating actions. Using a segmentation mask for selecting interesting objects is similarly adopted by prior works [33, 59]. Finally, this design leads to interpretable visual representations, which is helpful for us to diagnose the visual sim-to-real gap.
@@ VI Task Implementation In Simulation
Figure 5: Task Decomposition as Privileged Information (for teacher only). Given a task, we decompose it into a sequential series of primitive subtasks. As shown above, the task of finding an object A, grasping it, and then dropping into an object B can be decomposed into a sequence of subtasks. In simulation, the subtask boundaries and transitions are well-defined, serving as a form of privileged information to the teacher. The teacher can leverage this information, and use a separate policy to learn each subtask. This expertise is then distilled into the student, which is conditioned on both visual camera feed and language commands.
As mentioned previously in Section IV, SLIM employs a task decompositional approach paired with progressive PEX for training a teacher policy to solve the long-horizon task. In this work, we use a multi-stage search and pick-and-place task as our exemplar long-horizon task (Fig. A.16). This task will also serve as a concrete example for illustrating the task decomposition. Given a language instruction (e.g.,Drop the blue cube into the green basket.”), the robot needs to locate an object A (the blue cube), pick it up, and then drop it into a container B (the green basket). This task can be naturally broken down into the following sequential steps, where object A and B are specified by the language command itself:
@@
    Search: The robot must first search for object A using the RGB images from the wrist-mounted ego-centric camera.
@@
    MoveTo: After locating object A, the robot should approach it until it is within the reach of the arm.
@@
    Grasp: The robot attempts to grasp and pick up object A.
@@
    SearchWithObj: After picking up object A, the robot should then search for container B. Within this stage and those after, there should be some coordination between the arm and locomotion movement to facilitate subtask success while avoiding dropping the grasped object.
@@
    MoveToWithObj: After locating container B, the robot should approach it until it is within the reach of the arm.
@@
    MoveGripperToWithObj: The robot then moves its gripper towards container B until reaching an appropriate position over the top of the container.
@@
    DropInto: Finally, the robot releases its gripper and drops object A into container B.
An illustration of the full task decomposition can be observed in Fig. 5. Note that in this figure, we append an auxiliary Idle subtask, which encourages the robot to adjust its arm back to its starting position after finishing DropInto. This serves as an optional subtask and is used purely to facilitate subsequent deployments. Herein after, we do not include Idle in evaluations or discussions as it is unrelated to the task.
Task decomposition offers many benefits. For one, we can implement reward functions for each subtask, aiding modularity and ease of implementation. For all subtasks, we incorporate a sparse subtask success reward. Additionally, for those requiring moving the robot (or gripper) to a target region (e.g., MoveTo, Grasp, etc.), we incorporate an additional dense delta distance-based shaping reward. For SearchWithObj and MoveToWithObj, we also incorporate an arm retract reward (another delta distance-based shaping reward), where the arm is encouraged to stay close to a predefined target pose, shown in Fig. 6. This serves to promote good vantage and discourage suboptimal behaviors after grasping such as looking at the ground when searching. Its impact will be investigated in Section VIII-B (No Arm Retract baseline).
Finally, task decomposition offers the extra advantage of allowing behavior priors to be flexibly integrated into the policy. Such behavior priors can significantly increase sim-to-real task success as we will show later in Section VIII-D. More details on the subtasks, rewards, and behavior priors are provided in Appendix A-A.
[…]
Category Technique Description Reference Importance
Rating (1-3)
[…]
a deterministic behavior that fails in the real world
[…]
mismatch in the absence of system identification
[…]
for better generalization in real
[…]